# Enhancing inflationary model predictions via refined slow-roll dynamics

### v3

@@ III.1 Starobinsky inflation
With the help of Eq. (28), one can easily solve the scalar field solution in terms of the e-folding number $N_{k},$ and the solutions of the slow-roll parameters at the leading order can be obtained as:
@@
and correspondingly, the bound on $n_{s}$ becomes:
@@
In fact, as can be seen, for $w_{\rm re}<1/3,$ it is obvious that there is an upper bound on $N_{k}$, which implies a bound on $n_{s}$ as well, i.e., $N_{k}\leq 56$ and $n_{s}<0.964.$ It also immediately puts a tight constraint on $N_{\rm re}$ and $T_{\rm re}$ (for $w_{\rm re}=0$) as
@@
However, during reheating, if we assume $w_{\rm re}>1/3,$ these constraints flips and in the $1\sigma$ domain, there arises a new bound with a lower value of $N_{k}$ as well as in $n_{s}$ as
@@
Considering these bounds into account, we plot $r$ as a function of $n_{s}$ in Fig. 3. Further, instead of looking in $1\sigma$ limit, if we consider the observational constraints proposed by the forthcoming experiments such as PRISM [94] and EUCLID [95], cosmic 21-cm surveys [67], and CORE experiments [68], which offer a precision enhancement of $10^{-3}$ in $n_{s}$, the constraints on $N_{k}$, $N_{\rm re}$ and $T_{\rm re}$ can be substantially improved [^1] The central value of $n_{s}$ is assumed to be $0.9672$ [35]. Then one can immediately refer from the Fig. 2 and 3 that for $w_{\rm re}<1/3,$ the scalar spectral index does not satisfy the constraints in $1\sigma$ level posed by future observations. Therefore, under these conditions, this model can be ruled out. The model can be survived if one assumes $w_{\rm re}>1/3,$ and the bound on $N_{\rm re}$ and $T_{\rm re}$ for the constraints posed by future observations becomes
@@ IV.1 Implementation of accurate inflationary dynamics through numerical methods
As mentioned earlier, this epoch of the slow-roll regime lasts as long as both of the slow-roll parameters are small, i.e., ($\epsilon_{1}\ll 1,~\epsilon_{2}\ll 1$). On the other hand, $\epsilon_{1}=1$ signals the end of inflation. It turns out that even before the end of inflation, these slow approximations break down. Despite this, we still rely on (leading-order) slow-roll approximations even at the end of inflation to determine the analytical solution of the background variables, assuming the contribution due to accurate dynamics in evaluating perturbed observables is negligible. However, it is crucial, as in this section, we shall learn the significance of obtaining the accurate values of not only the background parameters but also the end of inflation. To illustrate this, let’s now discuss this in detail with an example: the Starobinsky inflation.
@@
along with the slow-roll solutions given in Eqs. (30). Here, $N_{\rm end}$ is the duration of the field rolling from the initial value $\phi_{i}$ to $\phi_{\rm end},$ i.e., the duration of inflation. These values are essential for the analysis in Eqs. (III) and (25) which has been shown in the previous section with Fig. 2.
@@
$H=\sqrt{\frac{V}{3-\frac{1}{2}\phi_{N}^{2}}},\quad\phi_{NN}+\left(3-\frac{1}{2}\phi_{N}^{2}\right)\left(\phi_{N}+\frac{V_{\phi}}{\phi}\right)=0.$
@@
Thus, there is a significant improvement not only in the numerical solution of the background variables but also in $H_{\rm end},$ $\rho_{\rm end},$ and $N_{\rm end}.$ These improvements directly impact the reheating analysis in two ways. The first is the direct influence of $\rho_{\rm end}$ in Eq. (III). The second, and more crucial, impact is the significant improvement in $\Delta N_{\rm end}=\Delta N_{k}\sim 1.5,$ which in turn enhances the accuracy of the background variables. For example, in the case of leading order approximations, $\{\epsilon_{1},\epsilon_{2}\}\simeq\{0.0003,0.040\}$ for $N_{k}=50,$ whereas, numerically we obtain these values as $\{\epsilon_{1},\epsilon_{2}\}\simeq\{0.00027,0.038\},$ and thus $\Delta n_{s}\simeq 2\times 10^{-3},$ as also discussed in Ref. [72]  [^2] Please note that, for a multi-parameter model, numerical evaluation for the entire range of the parameters can be challenging. In that case, one can rely upon a semi-analytical solution presented in Ref. [72], where analytically obtained solutions are nearly comparable to the numerical solutions..
@@
We now examine the impact of this on the reheating analysis and compare the results with the results obtained using slow-roll approximations. This is illustrated in Fig. 4. As anticipated, the improvement is significant. In the figure, one can see that numerical predictions show a significant shift to the right in the estimates of $N_{\rm re}$ and $T_{\rm re}.$ Due to this shift, the bound on $N_{k}$, and subsequently $n_{s}$ also changes, with an improvement being $\Delta n_{s}\sim 10^{-3}$. For example, in the case of leading-order slow-roll approximation with $w_{\rm re}<1/3,$ the upper bound on $n_{s}$ is: $n_{s}\leq 0.964,$ whereas, in this improved case, the bound becomes $n_{s}\leq 0.965.$ The improvements are explicitly shown in the Table 1. This confirms our anticipation that the accurate dynamics help improve the theoretical predictions, which is one of the main results of this work.
@@ IV.2 Implementation of higher-order slow-roll approximations
where $C=-2+\ln{2}+\gamma$, where $\gamma$ is the Euler constant, $f=5$ and $g=7$. In this section, instead of using Eqs. (16), along with the numerical solution, we consider the above expressions for the observable, repeat the reheating analysis, and compare our results with the numerical results obtained in the previous section, which has been illustrated in Figs. 6 and 7 as well as in Table 2.
@@
As again can be seen in Fig. 6, there is a further shift in reheating parameters, $N_{\rm re}$ and $T_{\rm re}$ compared to the improvements mentioned in the previous section. We obtain an additional improvement of $\Delta n_{s}\sim 4\times 10^{-4},$ resulting a new bound on $n_{s}$ becomes $n_{s}\leq 0.9653$ for $w_{\rm re}<1/3.$ Thus, considering higher-order slow-roll approximations instead of the leading-order can significantly improve the estimation of the observables by an order of $\sim 10^{-3}$. The detailed bounds on parameters are shown in the Table 2.
@@ IV.3 Implementation of the onset of reheating as the bottom of the potential
Till now, we were estimating $N_{\rm re}$ and $T_{\rm re}$, considering the onset of the reheating era after the end of inflation. However, in Ref. [63], it has been shown that, instead of considering minimal gravity theory, if one considers modified gravity theories, the onset of reheating as the end of inflation can a make a significant change in the analysis. This is because, $\epsilon_{1}$ is not a invariant under conformal transformation. Therefore, in a meaningful manner, one can choose the onset of reheating as the bottom of the potential, as this point is conformally invariant, and can be used even for general modified gravity theories. Therefore, in this section, we introduce corrections to the reheating analysis, and thus the observables, by considering that the reheating era commences after the epoch at which the potential touches its minima (bottom of the potential). In that case, Eq. (III) can be re-written as
$N_{\text{r}e}$ $=$ $\frac{4}{1-3w_{\text{re}}}\left(-N_{k}-N_{\text{eb}}-\frac{1}{4}\log\left(\frac{30}{\pi^{2}g_{\text{reh}}}\right)-\frac{1}{3}\log\left(\frac{11g_{\text{s,re}}}{43}\right)-\log\left(\frac{k}{a_{0}T_{0}}\right)\right$. (41)
$\left.-\frac{1}{4}\log\rho_{\text{b}}+\log H_{k}\right),$
@@
Here, $N_{\rm eb}\equiv a_{\rm b}/a_{\rm end}$ and $\rho_{\rm b}\equiv 3H_{\rm b}^{2},$ where $(b)$ denotes the bottom of the potential. Again, we perform our analysis considering the numerically obtained solution with higher-order slow-roll corrections and the onset of reheating as the bottom of the potential. However, to our surprise, we find that this implementation does not improve the accuracy to a significant amount, as $\Delta n_{s}\sim 10^{-5}.$
Figure 8: We plot the duration of reheating $N_{\rm re}$ and reheating temperature $T_{\rm re}$ given by Eqs. (III) and (25) as function of the scalar spectral index $n_{s}$ parametrically comparing analytically approximated solution and numerical solution with higher-order approximations, given by Eqs. (31), (39) respectively. The solid lines are for the onset of reheating as the bottom of the potential and the dashed lines are for the analytically approximated solution. Please note that different colors represent dynamics corresponding to different effective equations of state parameter $w_{\rm re}$ as indicated in the figure. The blue-shaded region represents the $1\sigma$ constraint on the value of $n_{s}$ using ongoing observations [32, 33, 34, 35] with $n_{s}=0.9672\pm 0.0059$. The dark blue region shows the future projected bound on $n_{s}$ with a sensitivity of $10^{-3}$, assuming its central value remains unchanged. The temperature below the lighter red region is excluded due to the constraint from the electroweak scale, which is taken to be 100100 GeV.
To summarize our cumulative result, in Fig. 8, we plot the results obtained through the analytical approximations mentioned in the first section and the improvement obtained through the implementation of three corrections to the dynamics, and Table 3 summarizes all the results. Compared to the conventional analytical case, our corrections improve the accuracy in $n_{s}$ to $\Delta n_{s}\sim 1.2\times 10^{-3},$ which is significant. Note that, if the central value of $n_{s}$ is assumed to be $\sim 0.9672$ even in this case, then even with the improvements, the Starobinsky model can indeed be ruled out at $1\sigma$ level from the expected future observation. These are the main results of this work.
@@ V Summary and conclusions
After implementing these changes, we found an improvement in the theoretical bounds on the observables. As seen in Fig. 8, there is an improvement of $\Delta n_{s}\sim 1.2\times 10^{-3}.$ These numerical predictions indicate that for $w_{\rm re}<1/3$, the Starobinsky model has an upper bound as $n_{s}\leq 0.9653.$ This can be ruled out from the future observational constraint if the central value is $n_{s}\simeq 0.9672$ [35]. However, for $w_{\rm re}>1/3$, the reheating parameters fall within the future observational bounds, though, at the cost of sufficient reheating e-folding number ($\sim\mathcal{O}(1-2)$). This scenario can again be ruled out if future observations detect instantaneous reheating [93, 92].
The result is compelling, as now it has been established that corrections due to accurate dynamics or higher-order slow-roll approximations can not be ignored. This also implies that even seemingly minor corrections must be considered and implemented to study their implications on perturbed observables. Therefore, when obtaining observational bounds on any inflationary models, one must consider all these corrections, which have been reserved for future endeavors.
In addition to these methods, other improvements can also be made to further enhance the accuracy. For example, in this work, we only considered the slow-roll dynamics, but one could also improve the reheating analysis by incorporating qualitative analysis of the reheating and implementing the changes into the quantitative analysis. Second, higher-order correlation such as non-Gaussianity (see, for instance, Refs. [96, 97, 98, 99, 100]) can play a pivotal role in the observables. Third, the implications of primordial black holes and primordial gravitational waves may also play a crucial role in inflationary and post-inflationary dynamics. Fourth, exploring the non-minimal sector instead of the minimal theory of gravity is another avenue. In fact, in our work, while the corrections due to the onset of reheating as the bottom of the potential did not play a significant role, they could be significant in the non-minimal sector [101, 63]. Finally, extending the analysis to non-inflationary dynamics, such as viable classical bounces, is also possible. Recent studies [102, 103, 100, 104, 105, 106, 107, 108] show that these alternatives can also fit observations well. Therefore, one must consider all alternatives to slow-roll inflation and carefully study the analysis with all effects, even those that seem minor, to determine the observable constraints of the model.
@@ Acknowledgements
DN is supported by the DST, Government of India through the DST-INSPIRE Faculty fellowship (04/2020/002142). MK is supported by a DST-INSPIRE Fellowship under the reference number: IF170808, DST, Government of India. DN, SY and MK are also very thankful to the Department of Physics and Astrophysics, University of Delhi. MK, SY and DN also acknowledge facilities provided by the IUCAA Centre for Astronomy Research and Development (ICARD), University of Delhi.

### v4

@@ III.1 Starobinsky inflation
With the help of Eq. (28), one can easily solve the scalar field solution in terms of the e-folding number $N_{k},$ and the solutions of the slow-roll parameters at the leading order, for $N_{k}\gg 1,$ can be obtained as:
@@
for $w_{\rm re}=0,$ and correspondingly, the bound on $n_{s}$ becomes:
@@
These bounds follow from imposing the observational constraint on $n_{s}$ together with the physical requirement $N_{\rm re}\geq 0$. In fact, as can be seen, for $w_{\rm re}<1/3,$ it is obvious that there is an upper bound on $N_{k}$, which implies a bound on $n_{s}$ as well, i.e., $N_{k}\leq 56$ and $n_{s}<0.964.$ It also immediately puts a tight constraint on $N_{\rm re}$ and $T_{\rm re}$ (for $w_{\rm re}=0$) as
@@
However, during reheating, if we assume $w_{\rm re}>1/3,$ these constraints flip and in the $1\sigma$ domain, there arises a new bound with a lower value of $N_{k}$ as well as in $n_{s}$ as
@@
Note that reheating temperatures below the BBN scale $(\sim 10~{\rm MeV})$ are observationally excluded; therefore, the region corresponding to $T_{\rm re}\leq 10^{-2}$ GeV is not physically viable and must be discarded. In addition, considerations related to the electroweak scale $(\sim 100~{\rm GeV})$ may motivate a higher reheating temperature in specific baryogenesis scenarios; however, this requirement is model-dependent and does not provide a constraint as robust as the BBN bound. Considering these bounds into account, we plot $r$ as a function of $n_{s}$ in Fig. 3. Further, instead of looking in $1\sigma$ limit, if we consider the observational constraints proposed by the forthcoming experiments such as PRISM [94] and EUCLID [95], cosmic 21-cm surveys [67], and CORE experiments [68], which offer a precision enhancement of $10^{-3}$ in $n_{s}$, the constraints on $N_{k}$, $N_{\rm re}$ and $T_{\rm re}$ can be substantially improved [^1] The central value of $n_{s}$ is assumed to be $0.9672$ [35]. Then one can immediately refer from the Fig. 2 and 3 that for $w_{\rm re}<1/3,$ the scalar spectral index does not satisfy the constraints in $1\sigma$ level posed by future observations. Therefore, under these conditions, this model would be disfavored at the $1\sigma$ level if the central value remains unchanged. The model can be survived if one assumes $w_{\rm re}>1/3,$ and the bound on $N_{\rm re}$ and $T_{\rm re}$ for the constraints posed by future observations becomes
@@ IV.1 Implementation of accurate inflationary dynamics through numerical methods
As mentioned earlier, this epoch of the slow-roll regime lasts as long as both of the slow-roll parameters are small, i.e., ($\epsilon_{1}\ll 1,~\epsilon_{2}\ll 1$). On the other hand, $\epsilon_{1}=1$ signals the end of inflation. It turns out that even before the end of inflation, these slow approximations break down. Despite this, we still rely on (leading-order) slow-roll approximations even at the end of inflation to determine the analytical solution of the background variables, assuming the contribution due to accurate dynamics in evaluating perturbed observables is negligible. While slow-roll remains an excellent approximation during most of the inflationary phase, its validity deteriorates near the end of inflation where $\epsilon_{1}\rightarrow 1$, and the higher-order kinetic contributions become non-negligible. However, it is crucial, as in this section, we shall learn the significance of obtaining the accurate values of not only the background parameters but also the end of inflation. To illustrate this, let’s now discuss this in detail with an example: the Starobinsky inflation.
@@
along with the slow-roll solutions given in Eqs. (30). In both the analytical and numerical cases, we impose the identical initial condition $\phi_{i}=5.8$ to ensure a consistent comparison. Here, $N_{\rm end}$ is the duration of the field rolling from the initial value $\phi_{i}$ to $\phi_{\rm end},$ i.e., the duration of inflation. These values are essential for the analysis in Eqs. (III) and (25) which has been shown in the previous section with Fig. 2.
@@
$H=\sqrt{\frac{V}{3-\frac{1}{2}\phi_{N}^{2}}},\quad\phi_{NN}+\left(3-\frac{1}{2}\phi_{N}^{2}\right)\left(\phi_{N}+\frac{V_{\phi}}{V}\right)=0.$
@@
Thus, the numerical treatment shifts the end of inflation by $\Delta N_{\rm end}\simeq 1.6$, which directly propagates into the reheating analysis. These improvements directly impact the reheating analysis in two ways. The first is the direct influence of $\rho_{\rm end}$ in Eq. (III). The second, and more crucial, impact is the significant improvement in $\Delta N_{\rm end}=\Delta N_{k}\sim 1.5,$ which in turn enhances the accuracy of the background variables. For example, in the case of leading order approximations, $\{\epsilon_{1},\epsilon_{2}\}\simeq\{0.0003,0.040\}$ for $N_{k}=50,$ whereas, numerically we obtain these values as $\{\epsilon_{1},\epsilon_{2}\}\simeq\{0.00027,0.038\},$ leading to a representative shift of $\Delta n_{s}\simeq 2\times 10^{-3}$ around $N_{k}\simeq 50,$ as also discussed in Ref. [72]  [^2] Please note that, for a multi-parameter model, numerical evaluation for the entire range of the parameters can be challenging. In that case, one can rely upon a semi-analytical solution presented in Ref. [72], where analytically obtained solutions are nearly comparable to the numerical solutions..
@@
We now examine the impact of this on the reheating analysis and compare the results with the results obtained using slow-roll approximations. This is illustrated in Fig. 4. As anticipated, the improvement is significant. In the figure, one can see that numerical predictions show a significant shift to the right in the estimates of $N_{\rm re}$ and $T_{\rm re}.$ Due to this shift, the bound on $N_{k}$, and subsequently $n_{s}$ also changes, with an improvement being $\Delta n_{s}\sim 10^{-3}$. For example, in the case of leading-order slow-roll approximation with $w_{\rm re}<1/3,$ the upper bound on $n_{s}$ is: $n_{s}\leq 0.964,$ whereas, in this improved case, the bound becomes $n_{s}\leq 0.965.$ The improvements are explicitly shown in Table 1. This confirms our anticipation that the accurate dynamics help improve the theoretical predictions, which is one of the main results of this work.
@@ IV.2 Implementation of higher-order slow-roll approximations
where $C=-2+\ln{2}+\gamma$, and $\gamma$ is the Euler constant, $f=5$ and $g=7$. Here, all slow-roll parameters are evaluated using the full numerical background solution. In this section, instead of using Eqs. (16), along with the numerical solution, we consider the above expressions for the observable, repeat the reheating analysis, and compare our results with the numerical results obtained in the previous section, which has been illustrated in Figs. 6 and 7 as well as in Table 2.
@@
As again can be seen in Fig. 6, there is a further shift in reheating parameters, $N_{\rm re}$ and $T_{\rm re}$ compared to the improvements mentioned in the previous section. We obtain an additional improvement of $\Delta n_{s}\sim 4\times 10^{-4},$ resulting a new bound on $n_{s}$, namely $n_{s}\leq 0.9653$ for $w_{\rm re}<1/3.$ Thus, considering higher-order slow-roll approximations instead of the leading-order leads to an additional correction of order $4\times 10^{-4}$, further refining the theoretical predictions. The detailed bounds on parameters are shown in the Table 2.
@@ IV.3 Implementation of the onset of reheating as the bottom of the potential
Till now, we were estimating $N_{\rm re}$ and $T_{\rm re}$, considering the onset of the reheating era after the end of inflation. However, in Ref. [63], it has been shown that, instead of considering minimal gravity theory, if one considers modified gravity theories, the onset of reheating as the end of inflation can make a significant change in the analysis. This is because, $\epsilon_{1}$ is not a invariant under conformal transformation. Therefore, in a meaningful manner, one can choose the onset of reheating as the bottom of the potential, as this point is conformally invariant, and can be used even for general modified gravity theories. Therefore, in this section, we introduce corrections to the reheating analysis, and thus the observables, by considering that the reheating era commences after the epoch at which the potential touches its minima (bottom of the potential). In that case, Eq. (III) can be re-written as
$N_{\text{r}e}$ $=$ $\frac{4}{1-3w_{\text{re}}}\left(-N_{k}-N_{\text{eb}}-\frac{1}{4}\ln\left(\frac{30}{\pi^{2}g_{\text{reh}}}\right)-\frac{1}{3}\ln\left(\frac{11g_{\text{s,re}}}{43}\right)-\ln\left(\frac{k}{a_{0}T_{0}}\right)\right$. (41)
$\left.-\frac{1}{4}\ln\left(\frac{\rho_{\text{b}}}{H_{k}^{4}}\right)\right),$
@@
Here, $N_{\rm eb}\equiv a_{\rm b}/a_{\rm end}$ and $\rho_{\rm b}\equiv 3H_{\rm b}^{2},$ where $(b)$ denotes the bottom of the potential. Again, we perform our analysis considering the numerically obtained solution with higher-order slow-roll corrections and the onset of reheating as the bottom of the potential. However, to our surprise, we find that this implementation does not improve the accuracy to a significant amount, as $\Delta n_{s}\sim 10^{-5}.$ This small correction reflects the fact that, in minimally coupled models, the difference between the end of inflation and the minimum of the potential is dynamically short.
Figure 8: We plot the duration of reheating $N_{\rm re}$ and reheating temperature $T_{\rm re}$ given by Eqs. (III) and (25) as function of the scalar spectral index $n_{s}$ parametrically comparing analytically approximated solution and numerical background solution with higher-order slow-roll corrections, given by Eqs. (31), (39) respectively. The solid lines are for the onset of reheating as the bottom of the potential and the dashed lines are for the analytically approximated solution. Please note that different colors represent dynamics corresponding to different effective equations of state parameter $w_{\rm re}$ as indicated in the figure. The blue-shaded region represents the $1\sigma$ constraint on the value of $n_{s}$ using ongoing observations [32, 33, 34, 35] with $n_{s}=0.9672\pm 0.0059$. The dark blue region shows the future projected bound on $n_{s}$ with a sensitivity of $10^{-3}$, assuming its central value remains unchanged. The temperature below the lighter red region is excluded due to the constraint from the electroweak scale, which is taken to be 100100 GeV.
To summarize our cumulative result, in Fig. 8, we plot the results obtained through the analytical approximations mentioned in the first section and the improvement obtained through the implementation of three corrections to the dynamics, and Table 3 summarizes the cumulative effect of all corrections relative to the purely analytical treatment. Compared to the purely analytical slow-roll treatment presented in Sec. III, the cumulative implementation of these corrections results in a maximum shift of $\Delta n_{s}\sim 1.2\times 10^{-3}$ within the allowed reheating range. This shift is evaluated relative to the leading-order slow-roll prediction at fixed reheating assumptions. Note that, if the central value of $n_{s}$ is assumed to be $\sim 0.9672$ even in this case, then even with the improvements, the Starobinsky model would be disfavored at the $1\sigma$ level from the expected future observation. Although illustrated using the Starobinsky model, similar end-of-inflation corrections are expected in any inflationary model where slow-roll breaks down near the end of inflation. These are the main results of this work.
@@ V Summary and conclusions
After implementing these changes, we found an improvement in the theoretical bounds on the observables. As seen in Fig. 8, there is an improvement of $\Delta n_{s}\sim 1.2\times 10^{-3}.$ These numerical predictions indicate that for $w_{\rm re}<1/3$, the Starobinsky model has an upper bound as $n_{s}\leq 0.9653.$ This can be disfavored from the future observational constraint if the central value is $n_{s}\simeq 0.9672$ [35]. However, for $w_{\rm re}>1/3$, the reheating parameters fall within the future observational bounds, though, at the cost of sufficient reheating e-folding number ($\sim\mathcal{O}(1-2)$). This scenario would be strongly constrained if future observations favor instantaneous reheating [93, 92].
These results demonstrate that an accurate theoretical treatment of end-of-inflation dynamics is as essential as higher-order perturbative corrections in precision-era inflationary analyses. Even corrections that are often regarded as subleading can induce observable shifts in perturbed quantities, and therefore must be systematically incorporated when deriving constraints on inflationary models. While the present work focuses on slow-roll dynamics within a minimal framework, further refinements — including a more detailed reheating treatment, higher-order correlations such as non-Gaussianity [96, 97, 98, 99, 100], the role of primordial black holes and gravitational waves, extensions to non-minimal gravity [101, 63], and alternative early-universe scenarios such as viable classical bounces [102, 103, 100, 104, 105, 106, 107, 108] — may further sharpen theoretical predictions. A careful and consistent implementation of such effects remains an important direction for future investigations.
@@ Acknowledgements
DN is supported by the DST, Government of India through the DST-INSPIRE Faculty fellowship (04/2020/002142). MK is supported by a DST-INSPIRE Fellowship under the reference number: IF170808, DST, Government of India. D.N. is grateful to the Department of Physics, School of Advanced Sciences, Vellore Institute of Technology (VIT) Chennai, for institutional support and academic encouragement. DN, SY and MK are also very thankful to the Department of Physics and Astrophysics, University of Delhi. MK, SY and DN also acknowledge facilities provided by the IUCAA Centre for Astronomy Research and Development (ICARD), University of Delhi.